Determine the equation of the line that passes through the point (8,2) and is parallel to the line 2 y - 16 x = 3.
Answer by scott8148(6628)

putting the equation in slope-intercept form gives y=8x+(3/2)

parallel means same slope, so using point-slope form for the new line gives y-2=8(x-8)

or y-8x=-62